I have to evaluate the function at the given value of the independent variable and simplify the following:
1. f(x)=-3x-8:f(-2)
2. f(x)=x^2-1:f(x-2)
3. f(x)= 4x^2+5x-6:f(x-1)
4. f(x)=x^2-6/x^3-7x :f(5)

Answer by Alan3354(69443)

evaluate the function at the given value of the independent variable and simplify the following:
1. f(x)=-3x-8:f(-2)
Sub -2 for x
f(-2) = -3*(-2) - 8 = 6-8
f(-2) = -2
--------------------------
Do the same for the others.
2. f(x)=x^2-1: f(x-2)
3. f(x)= 4x^2+5x-6: f(x-1)
4. f(x)=x^2-6/x^3-7x: f(5)